OpenStudy (anonymous):

Chris is comparing two sales positions that he has been offered. The first pays a weekly salary of $375 plus a 20% commission. The second pays a weekly salary of $325 plus a 25% commission. How much must he make in sales per week for the two jobs to pay the same?

OpenStudy (anonymous):

375+0.2x=325+0.25x 0.05x=375-325 x=1000 $ each one must sale 1000 $ to make both of them take the same salary.